# SWE-Compass: Towards Unified Evaluation of Agentic Coding Abilities for Large Language Models

### v2

@@ 1 Introduction
  • A comprehensive, execution-grounded benchmark for software engineering.: We introduce SWE-Compass, a large-scale benchmark comprising 2,000 curated instances that span eight task types, eight programming scenarios, and ten programming languages. Each instance is drawn from real-world GitHub pull requests and paired with a reproducible execution environment, enabling rigorous and faithful evaluation of model performance in realistic development workflows.
  • A systematic evaluation framework aligned with real-world developer activities.: SWE-Compass establishes a structured taxonomy to assess models across different dimensions such as feature implementation, refactoring, test generation, and deployment. This design enables fine-grained diagnosis of LLM capabilities and provides a principled foundation for comparing agentic coding systems under consistent conditions.
  • Comprehensive empirical analysis and insights into LLM coding behavior.: Experiments with state-of-the-art LLMs and agentic systems reveal persistent gaps across tasks, languages, and scenarios, highlighting the difficulty of scaling beyond bug fixing and emphasizing the need for benchmarks that reflect the full complexity of real-world software engineering.
@@ 3.4 Evaluation Metrics
    LLM-As-A-Judge Score: This method (e.g., (Zheng et al., 2023b; Zhang et al., 2025b; c; Li et al., 2025)) uses a large language model (LLM) to review the model output according to a checklist; the final score is the proportion of checkpoints passed by the model output.
For specific tasks, the following metrics are used: Feature Implementation, Feature Enhancement, Bug Fixing, and Refactoring: For these tasks, Pass@1 is used to measure the model’s performance. Performance Optimization: For this task, the Performance Optimization Score is used to evaluate the model’s performance. Test Case Generation: We employ Line Coverage to assess the quality of the test cases generated by the model. In our implementation, we use pytest  (Pajankar, 2017) to compute line coverage for Python. For TypeScript and JavaScript, we use C8 (Vassudanagunta, 2025) to calculate line coverage. Code Understanding: For this task, we use the LLM-As-A-Judge Score to evaluate the accuracy of the model’s understanding of the code. The specific prompt can be found in the Appendix A.1.
@@ 4.2.2 Further Analysis
Figure 4: Comparison of Pass@1 (%) across the top programming languages for the two agent workflows (SWE-Agent vs Claude Code). Bars represent Pass@1, and overlaid markers indicate per-model average interaction turns; languages are ordered by overall Pass@1. This plot highlights whether improvements are concentrated in specific languages.
Figure 5: Distribution of interaction turns required per language for selected models, with a secondary axis or overlay showing Pass@1 to reveal trade-offs between effort (turns) and success. This highlights whether models achieve high Pass@1 by spending more turns on particular languages.
@@
Figure 6: (Two panels) (a) Scatter showing per-model aggregate Pass@1 vs per-language median Pass@1 with fitted line and Spearman $\rho$ to measure rank consistency; (b) Scatter of aggregate mean performance vs per-model performance variance (std) across languages/tasks to assess specialization. Together, these panels reveal if high overall scores are driven by gains on a single language (e.g., Python) and quantify consistency.
@@
We also retain a small residual bucket OTH (Other) for rare cases not covered by the above taxonomy (e.g., corrupted artifacts or external executor glitches). The figure caption lists all abbreviations for completeness. As shown in Figure 7, we report the per-model distribution of failure modes on SWE-Compass. Based on 600 error traces per model, we draw the following conclusions: (1) Shared bottlenecks in comprehension and implementation. All models exhibit high error rates in Requirement Misinterpretation (30–34%) and Incomplete Solution & Side Effects (29–42%), together accounting for $60\%$ of failures. By contrast, Technical Knowledge Gap is consistently low (5–8%), suggesting the core limitations lie in requirement grounding and holistic solution design rather than basic coding proficiency. (2) Distinct model characteristics. Claude-Sonnet-4 is the most balanced, showing the lowest Technical Knowledge Gap (4.7%) but room to improve on Inadequate Testing (20.8%). Qwen3-Coder-480B has the highest Incomplete Solution & Side Effects rate (42%, vs. Claude-Sonnet-4’s 32.7%), revealing weaknesses in end-to-end design. Gemini-2.5-Pro shows the highest Requirement Misinterpretation (34%) and a notable Infinite Loop issue (8.3%), posing reliability risks in production.
@@ 5 Conclusion
We introduced SWE-Compass, a unified, execution-grounded benchmark that enables systematic evaluation of large language models across diverse software engineering tasks, scenarios, and languages. By integrating 2,000 verified instances derived from real-world GitHub repositories with reproducible execution environments, SWE-Compass provides comprehensive coverage of the software development lifecycle. Our large-scale experiments with ten state-of-the-art LLMs under two agentic frameworks reveal consistent hierarchies of task difficulty, language-specific variability, and dominant failure modes rooted in requirement misinterpretation and incomplete solutions. These findings highlight that future progress in automated software engineering depends less on isolated code generation improvements and more on enhancing requirement grounding, environment reliability, and reasoning consistency. SWE-Compass offers a rigorous, scalable, and reproducible foundation for advancing the next generation of robust, general-purpose coding agents.
@@ A.4 Executor Details and Method-Specific Settings
Figure 8: Distribution of modified files and lines in golden patches, aggregated by benchmark slices and task types. The breakdown shows the scope and granularity of code changes across different programming tasks.

### v3

@@ 1 Introduction
  • A comprehensive, execution-grounded benchmark for software engineering. We introduce SWE-Compass, a large-scale benchmark comprising 2,000 curated instances that span eight task types, eight programming scenarios, and ten programming languages. Each instance is drawn from real-world GitHub pull requests and paired with a reproducible execution environment, enabling rigorous and faithful evaluation of model performance in realistic development workflows.
  • A systematic evaluation framework aligned with real-world developer activities. SWE-Compass establishes a structured taxonomy to assess models across different dimensions such as feature implementation, refactoring, test generation, and deployment. This design enables fine-grained diagnosis of LLM capabilities and provides a principled foundation for comparing agentic coding systems under consistent conditions.
  • Comprehensive empirical analysis and insights into LLM coding behavior. Experiments with state-of-the-art LLMs and agentic systems reveal persistent gaps across tasks, languages, and scenarios, highlighting the difficulty of scaling beyond bug fixing and emphasizing the need for benchmarks that reflect the full complexity of real-world software engineering.
@@ 3.4 Evaluation Metrics
    LLM-As-A-Judge Score: Following (Zheng et al., 2023b; Zhang et al., 2025b; c; Li et al., 2025)), we use a large language model (LLM) to review the model output according to a checklist; the final score is the proportion of checkpoints passed by the model output.
For specific tasks, the following metrics are used. For Feature Implementation, Feature Enhancement, Bug Fixing, and Refactoring, Pass@1 is used to measure the model’s performance. For Performance Optimization, the Performance Optimization Score is used to evaluate the model’s performance. For Test Case Generation, we employ Line Coverage to assess the quality of the test cases generated by the model. In our implementation, we use pytest  (Pajankar, 2017) to compute line coverage for Python. For TypeScript and JavaScript, we use C8 (Vassudanagunta, 2025) to calculate line coverage. For Code Understanding, we use the LLM-As-A-Judge Score to evaluate the accuracy of the model’s understanding of the code. The specific prompt can be found in the Appendix A.1.
@@ 4.2.2 Further Analysis
Figure 4: Comparison of Pass@1 (%) across the top programming languages for SWE-Agent. Bars represent Pass@1; languages are ordered by overall Pass@1. This plot highlights whether improvements are concentrated in specific languages.
Figure 5: Distribution of interaction turns required per language for selected models to reveal trade-offs between effort (turns) and success. This highlights whether models achieve high Pass@1 by spending more turns on particular languages.
@@
Figure 6: (Two panels)(a) This plot illustrates the relationship between skill specialization and performance consistency. Skill specialization is measured by the variability in performance across different tasks/languages, with higher values indicating stronger performance in specific tasks/languages compared to others. Performance consistency, on the other hand, reflects how stable the performance is across different difficulty levels, with higher values indicating more consistent performance over time. (b) The graph shows the relationship between performance variability and overall performance. Performance variability measures how much performance fluctuates across different tasks/languages, with higher values indicating greater inconsistency. Overall performance is calculated as the average score across tasks/languages, with higher values indicating better overall performance.
@@
Note that OTH (Other) denotes rare cases, which are not covered by the above taxonomy (e.g., corrupted artifacts or external executor glitches). The figure caption lists all abbreviations for completeness. As shown in Figure 7, we report the per-model distribution of failure modes on SWE-Compass. Based on 600 error traces per model, we draw the following conclusions: (1) Shared bottlenecks in comprehension and implementation. All models exhibit high error rates in Requirement Misinterpretation (30–34%) and Incomplete Solution & Side Effects (29–42%), together accounting for $60\%$ of failures. By contrast, Technical Knowledge Gap is consistently low (5–8%), suggesting the core limitations lie in requirement grounding and holistic solution design rather than basic coding proficiency. (2) Distinct model characteristics. Claude-Sonnet-4 is the most balanced, showing the lowest Technical Knowledge Gap (4.7%) but room to improve on Inadequate Testing (20.8%). Qwen3-Coder-480B has the highest Incomplete Solution & Side Effects rate (42%, vs. Claude-Sonnet-4’s 32.7%), revealing weaknesses in end-to-end design. Gemini-2.5-Pro shows the highest Requirement Misinterpretation (34%) and a notable Infinite Loop issue (8.3%), posing reliability risks in production.
@@ 5 Conclusion
We introduced SWE-Compass, a unified benchmark that enables systematic evaluation of large language models across diverse software engineering tasks, scenarios, and languages. By integrating 2,000 verified instances derived from real-world GitHub repositories with reproducible execution environments, SWE-Compass provides comprehensive coverage of the software development lifecycle. Our large-scale experiments with ten state-of-the-art LLMs under two agentic frameworks reveal consistent hierarchies of task difficulty, language-specific variability, and dominant failure modes rooted in requirement misinterpretation and incomplete solutions. These findings highlight that future progress in automated software engineering depends less on isolated code generation improvements and more on enhancing requirement grounding, environment reliability, and reasoning consistency. SWE-Compass offers a rigorous, scalable, and reproducible foundation for advancing the next generation of robust, general-purpose coding agents.
@@ A.4 Executor Details and Method-Specific Settings
Figure 8: Distribution of modified files and lines involved in golden patches.